OpenStudy (blaque678):

Find the solution of the exponential equation, correct to four decimal places. e^-4x=8 x=

OpenStudy (anonymous):

lake the natural logarithm of both sides: \[\ln(e^{-4x}) = \ln(8)\] since the natural log is the inverse operation of the exponential function: \[-4x = \ln(8)\] \[x = \frac{ -\ln(8) }{ 4 }\]

OpenStudy (blaque678):

4 decimal places.....

OpenStudy (anonymous):

use your calculator....